# Bounding the degree of generic sharp transitivity

### v1

@@ 2.3. Groups of rank 6
That $G$ has no rank $5$ subgroups is due to Hrushovski, see [BN94, Theorem 11.98].
Suppose $M<G$ is nonsolvable and connected of rank $4$. Our general strategy is as follows: (1) use the presence of the nonsolvable $M$ of rank $4$ to produce a well-structured solvable $D$ of rank $4$, (2) study $(G,G/D)$ to identify $G$ as $\operatorname{PSL}_{2}(F)$, and (3) note that $\operatorname{PSL}_{2}(F)$ has no subgroup like $M$.
@@
Consider $N:=N_{G}^{\circ}(W)$. It contains $\langle G_{1,2}^{\circ},F^{\circ}(G_{1}^{\circ}),F^{\circ}(G_{2}^{\circ})\rangle$, so $\operatorname{rk}N$ is at least $4$, hence equal to $4$. If $N$ is nonsolvable, the analysis of Claim 2 could be applied to $N$ to force $W$ to be a torus, a contradiction. So $N$ is solvable, and by Claim 1, $N$ is nonilpotent. Thus $F:=F^{\circ}(N)$ has rank $2$ or $3$.
@@
Next suppose $\operatorname{rk}F=3$. Since $S$ does not commute with $F^{\circ}(G_{1}^{\circ})$, $S$ is not contained in $F$, so $F=\langle F^{\circ}(G_{1}^{\circ}),F^{\circ}(G_{2}^{\circ})\rangle$. Then $F^{\circ}(G_{1}^{\circ})$ is corank $1$ in the nilpotent group $F$, so $F$ normalizes $F^{\circ}(G_{1}^{\circ})$. Of course $F^{\circ}(G_{1}^{\circ})$ is normal in $G_{1}^{\circ}$, and by rank considerations, $G_{1}^{\circ}$ is the normalizer in $G$ of $F^{\circ}(G_{1}^{\circ})$. Thus, $F\leq G_{1}^{\circ}$. However, we know the structure of $F$ (unipotent of rank $3$) and of $G_{1}^{\circ}$ (rank $4$ containing a rank $2$ good torus), so we have a contradiction.
@@
We appeal to theory of split $2$-transitive permutation groups (via the theory of Moufang sets). Since both factors in the splitting $G_{1}=F^{\circ}(G_{1})\rtimes G_{12}$ are abelian, we find that $G\cong\operatorname{PSL}_{2}(F)$ (see for example [Wis10, Corollary 1.2]). However, proper definable subgroups of $\operatorname{PSL}_{2}(F)$ are solvable while $M$ is not. ∎
Corollary 2.11.
If $G$ is a quasisimple group of Morley rank $6$ with finite center, then $G$ is an $N_{\circ}^{\circ}$-group.
@@
This follows immediately from Proposition 3 together with the fact that connected groups of rank at most $2$ are solvable. ∎
3. Proof of the main result
Throughout this final section, we adopt the hypotheses of the Theorem as well as the additional notation presented in Section 1.3. In particular, $(G,X)$ is a generically sharply $t$-transitive permutation group of finite Morley rank and $r=\operatorname{rk}X$. By existing work of Hrushovski (see [BN94, Theorem 11.98]) and the present authors [AW18], we may assume $r\geq 3$.
As indicated in Section 1.3, Proposition 1 will reduce our work to consideration of two case: (1) $G_{[t-1]}$ is solvable or (2) $r=3$ with $G_{[t-1]}\cong\operatorname{PGL}_{2}(F)$. We treat these two cases first and then our final proof of the Theorem will simply tie things together. We begin with the solvable case—this falls quickly thanks to Proposition 2.
@@ 3. Proof of the main result
Assume $G_{[t-1]}$ is solvable and, towards a contradiction, that $t\geq r+3$. Then $\Sigma_{t}\cong\operatorname{Sym}(t-1)\geq\operatorname{Sym}(r+2)$, and we consider the faithful action of $\Sigma_{t}$ on $G_{[t-1]}$ (see Section 1.3).
By [AW24, Lemma 2.7], $G_{[t-1]}$ is abelian, and then [CDW23] takes over—via the First Geometrisation Lemma followed by the Recognition Lemma—to show that $G_{[t-1]}$ is an elementary abelian $p$-group for some $p$ dividing $t-1$. However, with $G_{[t-1]}$ is abelian, Proposition 2 applies to the action of $G_{[t-2]}$ on its generic orbit, and we find that $G_{[t-1]}$ has a quotient isomorphic to $L^{\times}$ for some algebraically closed field $L$, a contradiction. ∎
@@
Assume $r=3$, $t\geq r+2$, and $G_{[t-1]}$ is nonsolvable. Then the following hold:
@@
Set $B=G_{[t-2]}$ and $H=G_{[t-1]}$.
@@
Let $G$ be an infinite simple algebraic group over a field $K$ of finite Morley rank. By the theorem of Macintyre, the field is algebraically closed, and thus, the structure theory of simple algebraic groups over algebraically closed fields applies. According to this, the dimension of $G$ over the pure field structure is $2u+l$ where $u$ is the dimension of the unipotent subgroup generated by a fixed set of positive roots and $l$ is the Lie rank of $G$ (see for example Section 28.5 of [Hum75]). Moreover, $u\geq l\geq 1$. It follows that $\operatorname{rk}(G)=(2u+l)\operatorname{rk}(K)\geq 3\operatorname{rk}(K)\geq 3$. If $l>1$ then the smallest value of $u$ is $3$ (see for example Section 9.3 of [Hum72]). Thus, $l=1$ since $\operatorname{rk}(G)\leq 6$. In this case, the only possibility for $G$ is $\operatorname{PSL}_{2}(K)$ (Corollary 32.3 of [Hum75]), and $\operatorname{rk}(K)$ is either $1$ or $2$. ∎
@@
Set $B=G_{[t-2]}$. Then $\operatorname{rk}(B)=6$, and $B$ is quasisimple with finite center according to Lemma 3.1(3). If $B$ contains a nontrivial unipotent $2$-subgroup, so does $B/Z(B)$. By the classification of the simple groups of finite Morley rank of even type ([ABC08]), $B/Z(B)$ is a simple linear algebraic group over an algebraically closed field of characteristic $2$. Lemma 3.2 then implies $B/Z(B)\cong\operatorname{PSL}_{2}(K)$ with $K$ algebraically closed. But $G_{[t-1]}$ is not solvable by Lemma 3.1(1), violating the structure of $\operatorname{PSL}_{2}(K)$. ∎
@@
Let $B=G_{[t-2]}$ and $C=C^{\circ}_{B}(i)$, and suppose $C$ is not solvable. By Proposition 3, $\operatorname{rk}C<4$, and the classification of nonsolvable groups of rank at most $3$ implies that $C\cong\operatorname{SL}_{2}(F)$ with $F$ of characteristic not $2$. But then the Sylow $2$-subgroup of $B$ matches that of $\operatorname{SL}_{2}(F)$, contrary to the fact that the Sylow $2$-subgroup of $G_{[t-1]}\leq B$ matches that of $\operatorname{PSL}_{2}(K)$ by Lemma 3.1(1). ∎
@@
We work with a counterexample and adopt the following assumptions: $r=3$, $t\geq r+2$, and $H=G_{[t-1]}$ is nonsolvable. Set $B=G_{[t-2]}$. By Lemma 3.1 and Corollaries 2.11 and 3.3, $B$ is an $N_{\circ}^{\circ}$-group of odd type, and by Lemma 3.4, we may apply Fact 2.10. The presence and structure of $H$ given in Lemma 3.1 implies that we are in type CiBo2.
@@
Next, notice that if $i\in H$, then $T:=(C\cap H)^{\circ}$ is a rank one good torus, which is the definable closure of a $2$-torus in $H$. In particular, $T$ is a characteristic rank $1$ subgroup of $C$. And as involutions are conjugate in $B$, $C^{\circ}(i)$ has such a characteristic rank $1$ subgroup, for every choice of $i\in I(B)$.
@@
Let $T$ be the definable closure of a $2$-torus in $C$. Then $S_{t-1,t}$ acts on the rank $1$ factors of the series $1\leq T\leq C$. Thus, $[\gamma,C]\leq T$, and $[\gamma,\gamma,C]=1$. If $[\gamma,C]\neq 1$, then (by usual quadratic arguments) $[\gamma,C]$ is an elementary abelian $3$-group, which also must be equal to $T$, a contradiction. We conclude that $[\gamma,C]=1$, so $C\leq C_{B}^{\circ}(\gamma)$.
By Lemma 3.1(1), so $\gamma$ acts on $H$ as an inner automorphism, $\gamma$ centralizes a rank $1$ connected subgroup $A$ of $H$. As observed above, $H\cap C=1$, so $\langle C,A\rangle$ is a connected subgroup of $C_{B}^{\circ}(\gamma)$ of rank at least $3$. Additionally, $\Sigma_{t-1,t}$ acts faithfully on $B$ (see Section 1.3), so $C_{B}^{\circ}(\gamma)$ is proper subgroup of $B$ that properly contains $C$, a contradiction. Thus, $G_{[t-1]}$ is solvable. ∎
@@
We work under the assumption that $t\geq r+2$. By the main theorem of [AW18], we may assume that $r\geq 3$, and using Proposition 4, we may also assume that $G_{[t-1]}$ is not solvable (aiming for a contradiction). Let $H=G_{[t-1]}$, and adopt the notation of Section 1.3.
Recall that $\operatorname{rk}(H)=\operatorname{rk}(X)=r$ and $\Sigma_{t}\cong\operatorname{Sym}(t-1)$ acts faithfully on $H$. When $t-1\geq 5$, Proposition 1 applies to this action—with one exception that we will treat below separately—and we have:
$\operatorname{rk}(H)\geq t-1\geq r+1.$
This is a contradiction to $\operatorname{rk}(H)=r$ unless we are in the exceptional case given in Proposition 1: $t-1=5$ and $H\cong\operatorname{PGL}_{2}(F)$ with $F$ algebraically closed. It follows from this that $r\in\{3,4\}$. Lemma 3.2 then implies that that $r=3$, and Proposition 5 gives the final contradiction. ∎

### v2

@@ 2.3. Groups of rank 6
Assume $G$ is a simple group of Morley rank $6$. That $G$ has no rank $5$ subgroups is due to Hrushovski, see [BN94, Theorem 11.98]. So we aim to show that connected rank $4$ subgroups are solvable. Once done, the final sentence of Proposition 3 then follows from the fact that connected groups of rank at most $2$ are solvable.
Now suppose $M<G$ is nonsolvable and connected of rank $4$. Our general strategy is as follows: (1) use the presence of the nonsolvable $M$ of rank $4$ to produce a well-structured solvable $D$ of rank $4$, (2) study $(G,G/D)$ to identify $G$ as $\operatorname{PSL}_{2}(F)$, and (3) note that $\operatorname{PSL}_{2}(F)$ has no subgroup like $M$.
@@
Consider $N:=N_{G}^{\circ}(W)$. It contains $\langle G_{1,2}^{\circ},F^{\circ}(G_{1}^{\circ}),F^{\circ}(G_{2}^{\circ})\rangle$, so $\operatorname{rk}N$ is at least $4$, hence equal to $4$. If $N$ is nonsolvable, the analysis of Claim 2 could be applied to $N$ to force $W$ to be a torus, a contradiction. So $N$ is solvable, and by Claim 1, $N$ is nonnilpotent. Thus $F:=F^{\circ}(N)$ has rank $2$ or $3$.
@@
Next suppose $\operatorname{rk}F=3$. Since $S$ does not commute with $F^{\circ}(G_{1}^{\circ})$, $S$ is not contained in $F$, so $F=\langle F^{\circ}(G_{1}^{\circ}),F^{\circ}(G_{2}^{\circ})\rangle$. Then $F^{\circ}(G_{1}^{\circ})$ is corank $1$ in the nilpotent group $F$, so $F$ normalizes $F^{\circ}(G_{1}^{\circ})$. Of course $F^{\circ}(G_{1}^{\circ})$ is normal in $G_{1}^{\circ}$, and by rank considerations, $G_{1}^{\circ}$ is the connected normalizer in $G$ of $F^{\circ}(G_{1}^{\circ})$. Thus, $F\leq G_{1}^{\circ}$. However, we know the structure of $F$ (unipotent of rank $3$) and of $G_{1}^{\circ}$ (rank $4$ containing a rank $2$ good torus), so we have a contradiction.
@@
We appeal to the theory of split $2$-transitive permutation groups (via the theory of Moufang sets). Since both factors in the splitting $G_{1}=F^{\circ}(G_{1})\rtimes G_{12}$ are abelian, we find that $G\cong\operatorname{PSL}_{2}(F)$ (see for example [Wis10, Corollary 1.2]). However, proper definable subgroups of $\operatorname{PSL}_{2}(F)$ are solvable while $M$ is not. ∎
This completes the proof of Proposition 3. ∎
3. Proof of the main result
Throughout this final section, we adopt the hypotheses of the Theorem as well as the additional notation presented in Section 1.3. In particular, $(G,X)$ is a generically sharply $t$-transitive permutation group of finite Morley rank, $(1,\ldots,t)\in X^{t}$ is a generic $t$-tuple, and $r=\operatorname{rk}X$. By existing work of Hrushovski (see [BN94, Theorem 11.98]) and the present authors [AW18], we may assume $r\geq 3$.
We begin by addressing connectedness of point-stabilizers, which we use in our analysis of $G_{[t-1]}$ and $G_{[t-2]}$.
Lemma 3.1.
If $t\geq 2$, then for every $0\leq m<t$, the pointwise stabilizer $G_{[m]}$ of $1,\ldots,m$ is connected of rank $(t-m)\cdot r$, where we take $G_{[0]}$ to be $G$.
@@ 3. Proof of the main result
Since $t\geq 2$, by [BC08, Lemma 1.8(3)], the set $X$ is connected; this is a key point.
Now, as the action of $G$ is generically sharply $t$-transitive, $G_{[m]}$ acts generically sharply $(t-m)$-transitively, so if $\mathcal{O}\subseteq X^{t-m}$ is the generic orbit of $G_{[m]}$, then $G_{[m]}$ is in definable bijection with $\mathcal{O}$ (making essential use of sharpness). Since $\mathcal{O}$ is generic in $X^{t-m}$, $\mathcal{O}$ has rank $(t-m)\cdot r$ as well as degree $1$ by our initial observation, so same is true of $G_{[m]}$. ∎
As indicated in Section 1.3, Proposition 1 will reduce our work to consideration of two cases: (1) $G_{[t-1]}$ is solvable or (2) $r=3$ with $G_{[t-1]}\cong\operatorname{PGL}_{2}(F)$. We treat these two cases first and then our eventual proof of the Theorem will simply tie things together. We begin with the solvable case—this falls quickly thanks to Proposition 2.
@@
Assume $G_{[t-1]}$ is solvable and, towards a contradiction, that $t\geq r+3$. Then $\Sigma_{t}\cong\operatorname{Sym}(t-1)\geq\operatorname{Sym}(r+2)$, and we consider the faithful action of $\Sigma_{t}$ on the rank $r$ group $G_{[t-1]}$ (see Section 1.3).
By [AW24, Lemma 2.7], $G_{[t-1]}$ is abelian, and then [CDW23] takes over—via the First Geometrisation Lemma followed by the Recognition Lemma—to show that $G_{[t-1]}$ is an elementary abelian $p$-group for some $p$ dividing $t-1$. However, since $G_{[t-1]}$ is abelian, Proposition 2 applies to the action of $G_{[t-2]}$ on its generic orbit, and we find that $G_{[t-1]}$ has a quotient isomorphic to $L^{\times}$ for some algebraically closed field $L$, a contradiction. ∎
@@
Assume $r=3$, $t\geq 5$, and $G_{[t-1]}$ is nonsolvable. Then the following hold:
@@
Set $B=G_{[t-2]}$ and $H=G_{[t-1]}$; recall from Lemma 3.1 that they are connected and of ranks $6$ and $3$ respectively.
@@
Our contradictory assumption that $G_{[t-1]}$ is nonsolvable forces $G_{[t-2]}$ to be quasisimple with a finite center and many involutions (due to the structure of $G_{[t-1]}$), and this creates serious tension with the Algebraicity Conjecture. Since $G_{[t-2]}$ has rank $6$, the conjecture implies—as seen in Lemma 3.3 below—that $G_{[t-2]}$ should be of the form $\operatorname{PSL}_{2}$ so should certainly not have a subgroup like $G_{[t-1]}$. This is exactly our contradiction when $G_{[t-2]}$ is of even type thanks to [ABC08]. In odd type, nonalgebraic $G_{[t-2]}$ remains a possibility, but we are able to sufficiently pin down its structure—this time thanks to Proposition 3 and [DJ16]—to again reach a contradiction.
@@
Let $G$ be an infinite simple algebraic group over a field $K$ of finite Morley rank. By a theorem of Macintyre, $K$ is algebraically closed, and thus, the structure theory of simple algebraic groups over algebraically closed fields applies. According to this, the dimension of $G$ over the pure field structure is $2u+l$ where $u$ is the dimension of the unipotent subgroup generated by a fixed set of positive roots and $l$ is the Lie rank of $G$ (see for example Section 28.5 of [Hum75]). Moreover, $u\geq l\geq 1$. It follows that $\operatorname{rk}(G)=(2u+l)\operatorname{rk}(K)\geq 3\operatorname{rk}(K)\geq 3$. If $l>1$ then the smallest value of $u$ is $3$ (see for example Section 9.3 of [Hum72]). Thus, $l=1$ since $\operatorname{rk}(G)\leq 6$. In this case, the only possibility for $G$ is $\operatorname{PSL}_{2}(K)$ (Corollary 32.3 of [Hum75]), and $\operatorname{rk}(K)$ is either $1$ or $2$. ∎
@@
Set $B=G_{[t-2]}$. Then $B$ is quasisimple with finite center according to Lemma 3.2(3). If $B$ contains a nontrivial unipotent $2$-subgroup, so does $B/Z(B)$. By the classification of the simple groups of finite Morley rank of even and mixed type ([ABC08]), $B/Z(B)$ is a simple linear algebraic group over an algebraically closed field of characteristic $2$. Lemma 3.3 then implies $B/Z(B)\cong\operatorname{PSL}_{2}(K)$ with $K$ algebraically closed, but $G_{[t-1]}$ is not solvable by Lemma 3.2(1), violating the structure of $\operatorname{PSL}_{2}(K)$. ∎
The tension in odd type remains. Things will begin to give as we invoke Proposition 3 and ultimately Fact 2.10 to show that if $G_{[t-2]}$ is not of the form $\operatorname{PSL}_{2}$ then it is still—in some sense—close (being of type CiBo2). In any case, the structure of $G_{[t-2]}$ becomes rather transparent and, with a little more effort, quickly falls apart.
The next lemma provides the bridge from Proposition 3 to Fact 2.10; the proof of Proposition 5 will follow.
@@
Let $B=G_{[t-2]}$ and $C=C^{\circ}_{B}(i)$. By [BC08, Proposition 5.15], $i\in C$. Suppose now that $C$ is not solvable. By Proposition 3, $\operatorname{rk}C\leq 3$; this is a crucial point that eliminates the possibility of $C$ containing a copy of $\operatorname{PSL}_{2}$.
The classification of nonsolvable groups of rank at most $3$ (and that $i\in C$) implies that $C\cong\operatorname{SL}_{2}(F)$ with $F$ of characteristic not $2$. But then the Sylow $2$-subgroup of $B$ matches that of $\operatorname{SL}_{2}(F)$, contrary to the fact that the Sylow $2$-subgroup of $G_{[t-1]}\leq B$ matches that of $\operatorname{PSL}_{2}(F)$ by Lemma 3.2(1). ∎
@@
We work with a counterexample; assume: $r=3$, $t\geq 5$, and $H=G_{[t-1]}$ is nonsolvable. Set $B=G_{[t-2]}$. By Lemma 3.2 and Corollary 3.4 together with Proposition 3, $B$ is an $N_{\circ}^{\circ}$-group of odd type, and by Lemma 3.5, we may apply Fact 2.10. The presence and structure of $H$ given in Lemma 3.2 implies that we are in type CiBo2.
@@
Next, notice that if $i\in H$, then $T:=(C\cap H)^{\circ}$ is a rank one good torus, which is the definable closure of a $2$-torus in $H$. In particular, $T$ is a characteristic rank $1$ subgroup of $C$. And as involutions are conjugate in $B$ (as a consequence of having type CiBo2), $C_{B}^{\circ}(i)$ has such a characteristic rank $1$ subgroup, for every choice of $i\in I(B)$.
@@
Let $T$ be the definable closure of a $2$-torus in $C$. Then $\Sigma_{t-1,t}$ acts on the rank $1$ factors of the series $1\leq T\leq C$. Thus, $[\gamma,C]\leq T$, and $[\gamma,\gamma,C]=1$. If $[\gamma,C]\neq 1$, then (by usual quadratic arguments) $[\gamma,C]$ is an elementary abelian $3$-group, which also must be equal to $T$, a contradiction. We conclude that $[\gamma,C]=1$, so $C\leq C_{B}^{\circ}(\gamma)$.
By Lemma 3.2(1), $\gamma$ acts on $H$ as an inner automorphism, so $\gamma$ centralizes a rank $1$ connected subgroup $A$ of $H$. As observed above, $H\cap C=1$, so $\langle C,A\rangle$ is a connected subgroup of $C_{B}^{\circ}(\gamma)$ of rank at least $3$. Additionally, $\Sigma_{t-1,t}$ acts faithfully on $B$ (see Section 1.3), so $C_{B}^{\circ}(\gamma)$ is proper subgroup of $B$ that properly contains $C$, a final contradiction. ∎
@@
We work towards a contradiction under the assumption that $t\geq r+3$. As we noted at the beginning of this section, we may assume that $r\geq 3$. Let $H=G_{[t-1]}$. By Lemma 3.1, $H$ is connected of Morley rank $r$.
By Proposition 4, $H$ is not solvable, and then invoking Proposition 5, we have that $r\geq 4$. We have seen that $H$ is connected of rank $r$ and $\Sigma_{t}\cong\operatorname{Sym}(t-1)$ acts faithfully on $H$. As $t-1\geq r+2\geq 6$, Proposition 1 implies that $t-1\leq\operatorname{rk}(H)=r$, a contradiction. ∎
Acknowledgements
We warmly thank the anonymous referee for a careful reading of the article and numerous helpful suggestions. Among other things, their remarks led us to clarify a number of points and to identify a small case that was overlooked in an earlier version of Proposition 1.